OpenStudy (anonymous):

How do you find the slope of a line? I can't find a way to learn it where it sinks in.

OpenStudy (anonymous):

rise over run

OpenStudy (anonymous):

Write your given equation of the line in the form of y=mx+b and compare that with y=mx+b the value of m will give you the slope.

OpenStudy (anonymous):

there are multiple ways to find it. if it gives you two points on the line what you do is use y^2-y^1/x^2-x^1 (not to the power of two or one it is the x and y cordinate numbers from the points) so if you were givin (3,1) and (8,4) you would write... 4-1/8-3 whitch equals 3/5 and it cannot be simplified so that is your slope. you can also do it y^1-y^2/x^1-x^2 but you have to do the same thing on the bottom for whatever you do on top. If it gives you an equation of a line in the form y=mx+b whatever takes the spot of m is the slope (b is the y intercept and it will always just stay as an x variable) so the slope of the line y=2/3x+4 the slope is 2/3. The equation of the line can be in multiple formats...

OpenStudy (anonymous):

@SofiaGambina

OpenStudy (anonymous):

it can say 7x+2y=8 then you would isolate the y, subtract 7 from both sides and then divide both sides by 2. y=-7/2x+4 so your slope would be -7/2

OpenStudy (anonymous):

lastly it may say a line that goes through a point is parallel to the line 1/2x+8 what is the slope? the lines are parallel so they have the same slope so the slope would be 1/2. you will learn about parallel and perpendicular lines later though so you dont really have to worry about this one or perpendicular. but perpendicular is the opposite reciprical so the slope for a perpendicular line to 1/2x+8 would be -2/1 or just -2

OpenStudy (anonymous):

@SofiaGambina